OpenStudy (anonymous):

A spring requires a force of 25 N to produce an extension of 2 cm in it. What is the work done in further extending it by 3 cm.

OpenStudy (anonymous):

applesbfhj

OpenStudy (rock_mit182):

\[F = k x\]The equation to find this : Where F = Spring's force x = displacement k = constant

OpenStudy (rock_mit182):

now, 1 newton(N) = \[Kg*\frac{ m }{ s ^{2} }\]

OpenStudy (rock_mit182):

so to have the same units in the equation you need to convert 2 cm to meters

OpenStudy (rock_mit182):

\[25 Kg \frac{ m }{ s ^{2} } = k * 2 cm(\frac{ 1m }{ 100cm})\]so you'll have: the displacemnt in meters cause the centimiters cancel each other out, now solve for k to find the constant, and finally multiply the value of the constant with the displacement 3cm Don't forget to convert 3 cm to meters